OpenStudy (lilmike234):

What is the root of the polynomial shown below? f(x)=x^3-x^2-9x+9

OpenStudy (anonymous):

is it multiple choice ?

OpenStudy (lilmike234):

Yes the answer choices are 9 -3 2 9

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=x^3-x^2-9x+9 }\) the first step is to factor the first two terms out of x², and to factor the last 2 terms out of -9. \(\large\color{black}{ \displaystyle f(x)=x^2(x-1)-9(x-1) }\)

OpenStudy (solomonzelman):

then, you know that (say I had): \(x(B)−s(B)\) then i would get \((x−s)(B)\) --------------------------------------- same principal is here (but your B is 'x-1') your x is x², and your s is 9

OpenStudy (lilmike234):

So my root would be 9 correct ?